OpenStudy (anonymous):

Write a proportion for the following problem. The tax on a new car costing $14,400 is $784. At the same tax rate, what is the tax on a vehicle costing $20,000? Round to the nearest dollar. 784/14,400 = x/20,000 784/14,400 = 20,000/x 20,000/14,400=784/x

OpenStudy (anonymous):

The answer is (A) 784/14,400 = x/20,000 show the works: 784/14,400 = x/20,000 20000 * 784 = 14400x x = 20000/14400 = x/784 x = 20000 * 784 /14400 x = $1089 The answer is (A) $1,089 Hope this help☻